OpenStudy (anonymous):

Evaluate the limit using L'Hospital's rule if necessary limx→∞ (x^4)e^(−x^2)

OpenStudy (anonymous):

\[\lim _{x \rightarrow \infty} x^4e ^{-x^2}\] That's what it looks like :)

OpenStudy (anonymous):

After simply plugging in 'infinity' for x, I get 'infinity/infinity' which is an indeterminate form and so I can use L'Hopital's Rule

OpenStudy (anonymous):

I'm not sure if I'm right though...

OpenStudy (anonymous):

And then I have trouble differentiating this xD

OpenStudy (irishboy123):

you re-wrote it as \(\frac{x^4}{e^{x^2}}\)?

OpenStudy (anonymous):

Yep! Exactly! Is that right?

OpenStudy (irishboy123):

do differentiate top and bottom...? what do you get?

OpenStudy (anonymous):

Is it \[\frac{ 4x^3 }{ 2xe ^{x^2} }\] ?

OpenStudy (irishboy123):

and cancel out x's

OpenStudy (anonymous):

Right! So... \[\frac{ 4x^2 }{ 2e ^{x^2} }\]

OpenStudy (anonymous):

But then it's still \[\frac{ \infty }{ }\]

OpenStudy (anonymous):

I mean infinity/ infinity

OpenStudy (irishboy123):

so do L'Hopital again

OpenStudy (irishboy123):

and you can factor the 4/2 as well :p not that it really matters

OpenStudy (anonymous):

No matter how many times I repeat won't it just keep giving me the same thing?

OpenStudy (anonymous):

Ahaha true! Kk :D

OpenStudy (anonymous):

Ohhh actually eventually I get 1/infinity right?!? :O

OpenStudy (irishboy123):

no. keep differentitaing the top and you'll get it down to a constant

OpenStudy (irishboy123):

yes \(e^{x^2}\) is a beast

OpenStudy (anonymous):

Haha awesomeness! And yeah it sure is a beast XD lol

OpenStudy (perl):

Correct, you will eventually get \( 1 / \infty \)

OpenStudy (anonymous):

And that's = 0 :D

OpenStudy (anonymous):

Ahhh I see! Yeah that makes sense! Thanks so much for the tips! :)